OpenStudy (bananas1234):

Solve for x. 3x2 – 12 = 0

OpenStudy (hockeychick23):

first step is to get 3x^2 by itself so we do: 3x^2-12+12=0+12 3x^2=12 do you know the next step?

OpenStudy (bananas1234):

no, sorry

OpenStudy (hockeychick23):

no problem, ok so we take 3x^2=12 and divide both sides by 3 3x^2/3=12/3 x^2=4 now we find the square root of x sqrt(x^2)= sqrt(4) x=2

OpenStudy (hockeychick23):

do you get how i got x=2?

OpenStudy (bananas1234):

i think so, you got the square root of 4?

OpenStudy (hockeychick23):

correct :) so when dividing 3x^2 by 3 you eliminate the constant (3) and are left with x^2 general rule is what you do to one side of the equation you also need to do to the other which is why i divided 12 by 3. once you divide 12/3 you are left with x^2=4 and by taking the square root we get x=2

OpenStudy (bananas1234):

i see

OpenStudy (bananas1234):

so whats next?

OpenStudy (hockeychick23):

finding that x=2 is your last step

OpenStudy (bananas1234):

ok

OpenStudy (bananas1234):

thank you!

OpenStudy (hockeychick23):

sure :)